A company uses the following standard costs to produce a single unit of output.

Direct materials 6 pounds at $1.20 per pound = $ 7.20
Direct labor 0.6 hour at $8.00 per hour = $ 4.80
Manufacturing overhead 0.6 hour at $4.00 per hour = $ 2.40



During the latest month, the company purchased and used 56,000 pounds of direct materials at a price of $1.50 per pound to produce 10,000 units of output. Direct labor costs for the month totaled $43,776 based on 5,760 direct labor hours worked. Variable manufacturing overhead costs incurred totaled $21,000 and fixed manufacturing overhead incurred was $18,000. Based on this information, the direct labor efficiency variance for the month was:

  • A. $4,224 favorable

  • B. $2,304 favorable

  • C. $1,920 unfavorable

  • D. $1,920 favorable

  • E. $2,304 unfavorable
Answer #1

Labor efficiency variance

= (SH - AH) *SR.

= (10,000*0.6 - 5760)*8

= (6000 - 5760)*8

= 1920 Favourable